# 2]Key Laboratory of System Control and Information Processing, Ministry of Education, Shanghai 200240, China

### v2

@@ Abstract
Inspired by the use of adaptive kernel-based Cohen’s class time-frequency distributions (CCTFDs) for cross-term suppression, this paper aims to explore novel adaptive kernel functions for denoising. We integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain. Some examples are also carried out to demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-arts in noise suppression.
@@ 1 Introduction
Cohen’s class time-frequency distribution (CCTFD) [1], also known as the bi-linear kernel function time-frequency distribution, is one of the most representative time-frequency analysis tools of the conventional time-frequency distributions. It includes particular cases the Wigner-Ville distribution (WVD) [2], the Choi-Williams distribution [3], the Kirkwood-Rihaczek distribution [4], the Born-Jordan distribution [5], the Zhao-Atlas-Marks distribution [6], the Margenau-Hill distribution [7], and the Page distribution [8]. Indeed, it can be regarded as a unified bi-linear time-frequency distribution and has found many applications in seismic exploration, electronic countermeasures, deep-sea detection, spectral imaging, and ultrasonic inspection [9, 10, 11, 12, 13, 14].
@@
The integral form of CCTFD can be written as the Fourier transform (FT) of the product of the ambiguity function and the kernel function. Cross-terms may arise from the bi-linear nature of CCTFD, causing interference and reducing clarity in signal analysis. There are a number of attempts to use the kernel of CCTFD to suppress cross-terms, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19]. However, these methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely. To overcome this shortcoming, Baraniuk et al. have proposed a series of methods using adaptive kernel functions to suppress cross-terms [20, 21, 22, 23]. The adaptive radial Gaussian kernel function-based CCTFD is one of the most famous in this field.
@@
The integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function. It is important to note that the kernel function mentioned here is a Fourier transform pair with the kernel function described above, as shown in Eq. (2). From the view point of signal processing, the CCTFD is a result of smoothing the WVD by using the kernel function. Inspired by the adaptive kernel function concept proposed by Baraniuk et al., we aim to implement adaptive denoising by designing an adaptive kernel function in the WVD domain. This approach will enable us to effectively extract the target signal from background noise and mitigate the impact of noise through the unique reconstruction property of WVD.
@@
An adaptive filter [24] is a type of filter that automatically adjusts its parameters in response to changes in the signal, making it particularly effective for noise reduction by adapting to varying characteristics of the noise and effectively suppressing it. The most celebrated result in this field is Wiener’s result [25] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method. The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [26, 27, 28, 29, 30, 31, 32].
@@
  • This paper proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
@@
The remainder of this paper is structured as follows. In Section 2, we recall some necessary background and notation on the convolution type of CCTFD. In Section 3, we provide the CCTFD-based adaptive filter method for additive noises jamming signals. In Section 4, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method. In Section 5, we draw a conclusion. All the technical proofs of our theoretical results are relegated to the appendix parts.
@@
$\mathrm{C}_{g}^{\mathrm{LSAF}}$ the least-squares adaptive filter-based CCTFD for the input signal $g$
@@ 2 Convolution type of CCTFD
The integral form of CCTFD of the function $f(\boldsymbol{x})\in L^{2}(\mathbb{R}^{N})$ reads [1, 33]
$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
@@ 3 Least-squares adaptive filter-based CCTFD
By using the orthogonal principle [34], the stationary assumption and the conventional convolution and correlation theorems to establish, simplify and solve the Wiener-Hopf equation, respectively, the least-squares adaptive filter transfer function in the WVD domain reads
@@
$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
@@ 4 Numerical experiments
This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain. It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs and the adaptive radial Gaussian kernel function-based CCTFD.
In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the 111-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively. For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively. Additionally, the filtering method utilizing the adaptive radial Gaussian kernel function-based CCTFD is referred to as the adaptive radial Gaussian kernel.
@@
In examples 1–4, additive white Gaussian noise is introduced to the signals, with the SNR ranging from $-5\mathrm{dB}$ to $5\mathrm{dB}$ and the observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$. The sampling frequencies are set as follows: $80\mathrm{Hz}$ for example 1, $100\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for example 3 and $175\mathrm{Hz}$ for example 4.
@@
Figures 1(a) and (b) plot respectively the SNR-MSE (logarithm base 101010) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using seven filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, Wiener filter and the proposed adaptive CCTFD. Figures 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these seven filtering methods. Figures 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these seven filtering methods. Figures 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these seven filtering methods. It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, and Wiener filters under different SNR levels.
@@
Figure 1: Performance evaluation of seven denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
@@
Figure 2: Performance evaluation of seven denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
@@
Figure 3: Performance evaluation of seven denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
@@
Figure 4: Performance evaluation of seven denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
@@
Additionally, we investigated the denoising performance of the seven methods under conditions of different colored noise. The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise. The observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4. Table 2 presents the MSE and PSNR values of the output signals through the seven filters under the influence of three types of colored noise. It can be seen that for blue noise, our proposed adaptive CCTFD significantly outperforms the other methods. For pink and red noise, our method is superior to the others for LFM and GELFM signals (except for the PSNR value of red noise in the GELFM signal, which is slightly lower than that of the Wiener filter). For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals. However, overall, our method consistently outperforms the fixed kernel function filtering methods and the adaptive radial Gaussian kernel function filtering method.
@@
Pink noise Margenau-Hill 0.0016-0.0016- 0.0016 0.1592-0.1592- 0.1592 0.0513-0.0513- 0.0513 0.1319-0.1319- 0.1319
Kirkwood-Rihaczek 0.06760.06760.0676 0.0050-0.0050- 0.0050 0.0330-0.0330- 0.0330 0.1200-0.1200- 0.1200
Born-Jordan 0.1582-0.1582- 0.1582 0.3053-0.3053- 0.3053 0.1123-0.1123- 0.1123 0.2134-0.2134- 0.2134
Page 0.09960.09960.0996 0.0360-0.0360- 0.0360 0.0201-0.0201- 0.0201 0.0582-0.0582- 0.0582
Adaptive radial Gaussian kernel 0.0451-0.0451- 0.0451 0.4237-0.4237- 0.4237 0.0662-0.0662- 0.0662 0.0510-0.0510- 0.0510
Wiener filter 0.4985-0.4985- 0.4985 0.6911-0.6911- 0.6911 $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ 0.3169-0.3169- 0.3169 0.3369-0.3369- 0.3369
Blue noise Margenau-Hill 0.14370.14370.1437 0.4563-0.4563- 0.4563 0.0637-0.0637- 0.0637 0.3730-0.3730- 0.3730
Kirkwood-Rihaczek 0.35800.35800.3580 0.3711-0.3711- 0.3711 0.00630.00630.0063 0.2404-0.2404- 0.2404
Born-Jordan 0.1057-0.1057- 0.1057 0.5265-0.5265- 0.5265 0.2016-0.2016- 0.2016 0.9358-0.9358- 0.9358
Page 0.32810.32810.3281 0.3355-0.3355- 0.3355 0.02390.02390.0239 0.06820.06820.0682
Adaptive radial Gaussian kernel 0.0481-0.0481- 0.0481 0.4892-0.4892- 0.4892 0.0671-0.0671- 0.0671 0.0191-0.0191- 0.0191
Wiener filter 0.4676-0.4676- 0.4676 0.6632-0.6632- 0.6632 0.5208-0.5208- 0.5208 1.0831-1.0831- 1.0831
@@
Red noise Margenau-Hill 0.06550.06550.0655 0.3004-0.3004- 0.3004 0.0624-0.0624- 0.0624 0.2233-0.2233- 0.2233
Kirkwood-Rihaczek 0.12430.12430.1243 0.2567-0.2567- 0.2567 0.0475-0.0475- 0.0475 0.2145-0.2145- 0.2145
Born-Jordan 0.0907-0.0907- 0.0907 0.3810-0.3810- 0.3810 0.1369-0.1369- 0.1369 0.3623-0.3623- 0.3623
Page 0.17200.17200.1720 0.2790-0.2790- 0.2790 0.0082-0.0082- 0.0082 0.0851-0.0851- 0.0851
Adaptive radial Gaussian kernel 0.01050.01050.0105 0.3455-0.3455- 0.3455 0.0743-0.0743- 0.0743 0.0395-0.0395- 0.0395
Wiener filter 0.8039-0.8039- 0.8039 1.2798-1.2798- 1.2798 $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ 0.4032-0.4032- 0.4032 0.3919-0.3919- 0.3919

### v3

@@ Abstract
Inspired by the use of adaptive kernel-based Cohen’s class time-frequency distributions (CCTFDs) for cross-term suppression, this paper aims to explore novel adaptive kernel functions for denoising. We integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain. Numerical experiments on both simulated and real-world signals demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-art methods in noise suppression.
@@ 1 Introduction
Cohen’s class time-frequency distribution (CCTFD) [Coh95], also known as the bi-linear kernel function time-frequency distribution, is one of the most representative time-frequency analysis tools of the conventional time-frequency distributions. It includes particular cases the Wigner-Ville distribution (WVD) [wigner1932quantum], the Choi-Williams distribution [zhao2021research], the Kirkwood-Rihaczek distribution [rihaczek1968signal], the Born-Jordan distribution [quan2021novel], the Zhao-Atlas-Marks distribution [zhao1990use], the Margenau-Hill distribution [liang2024spatial], and the Page distribution [xia1996generalized]. Indeed, it can be regarded as a unified bi-linear time-frequency distribution and has found many applications in seismic exploration, electronic countermeasures, deep-sea detection, spectral imaging, and ultrasonic inspection [brynolfsson2021time, wang2021matching, xu2023dtftcnet, vitor2023induction, han2021passive, ding2023regional]. In addition to traditional CCTFD and their variants, deep learning-based adaptive time-frequency analysis methods have attracted attention for their ability to learn optimal representations from data. These methods are useful for signal filtering, denoising, and separation, and have shown effectiveness in tasks such as video, speech, and radar signal analysis [guo2023deep, pan2022tfa, li2023incipient, samavat2022deep, hamdaoui2023improved, zhang2024adaptive].
@@
The integral form of CCTFD can be written as the Fourier transform (FT) of the product of the ambiguity function and the kernel function. Cross-terms may arise from the bi-linear nature of CCTFD, causing interference and reducing clarity in signal analysis. There are a number of attempts to use the kernel of CCTFD to suppress cross-terms, giving birth to many useful and effective kernel functions of which the CCTFD works well [qu2020radar, yao2012reducing, lopac2021detection, jadhav2022automated, li2023time]. However, these methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely. To overcome this shortcoming, Baraniuk et al. have proposed a series of methods using adaptive kernel functions to suppress cross-terms [baraniuk1991radially, 212733, 258128, jones1995adaptive]. The adaptive radial Gaussian kernel function-based CCTFD is one of the most famous in this field.
However, existing adaptive kernel methods have been primarily developed with the objective of cross-term suppression rather than signal denoising. These approaches adapt the kernel shape based on the local structure of the ambiguity function to enhance time-frequency concentration and reduce interference between components. While such techniques improve the readability of time-frequency representations, they do not explicitly address the problem of noise suppression or signal reconstruction optimality. Consequently, when the signal is heavily contaminated by noise, conventional adaptive kernel methods often fail to achieve the desired filtering performance.
@@
To overcome this limitation, we reformulate the adaptive kernel design problem from the perspective of optimal time-frequency filtering. The integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function. It is important to note that the kernel function mentioned here is a Fourier transform pair with the kernel function described above, as shown in Eq. (2). From the viewpoint of signal processing, the CCTFD is a result of smoothing the WVD by using the kernel function. Inspired by the adaptive kernel function concept proposed by Baraniuk et al., we aim to implement adaptive denoising by designing an adaptive kernel function in the WVD domain. This approach will enable us to effectively extract the target signal from background noise and mitigate the impact of noise through the unique reconstruction property of WVD.
@@
An adaptive filter [diniz1997adaptive] is a type of filter that automatically adjusts its parameters in response to changes in the signal, making it particularly effective for noise reduction by adapting to varying characteristics of the noise and effectively suppressing it. The most celebrated result in this field is Wiener’s result [wiener1949extrapolation] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method. The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [ceccato2020spatial, baudais2023doppler, plabst2020wiener, song2020two, chang2022tensor, qi2021wiener, vered2024parallel]. Unlike previous adaptive kernel designs focused on cross-term suppression, the proposed method establishes a framework that bridges time-frequency representation and adaptive filtering theory, thereby providing a fundamental advance in both interpretability and performance.
@@
  • This paper proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extremely strong additive noise background.
@@
The remainder of this paper is structured as follows. In Section 2, we recall some necessary background and notation on the convolution type of CCTFD. In Section 3, we provide the CCTFD-based adaptive filter method for additive noises jamming signals. In Section 4, we introduce numerical examples and real-world data experiments to validate the effectiveness, reliability, and feasibility of the proposed method. In Section 5, we draw a conclusion. All the technical proofs of our theoretical results are relegated to the appendix parts.
@@
$\mathrm{C}_{g}^{\mathrm{LSAF}}$ the least-squares adaptive filter-based CCTFD
for the input signal $g$
@@ 2 Convolution type of CCTFD
The integral form of CCTFD of the function $f(\boldsymbol{x})\in L^{2}(\mathbb{R}^{N})$ reads [Coh95, ozaktas1996effect]
$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
$\times\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
@@ 3 Least-squares adaptive filter-based CCTFD
By using the orthogonal principle [proakis1985probability], the stationary assumption and the conventional convolution and correlation theorems to establish, simplify and solve the Wiener-Hopf equation, respectively, the least-squares adaptive filter transfer function in the WVD domain reads
@@
$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
$\times\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
@@ 4 Numerical experiments
This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain. The experiments include both simulated signals and real-world data from the Radar Characterisation (RadChar) dataset [Huang23Multi]. The denoising performance of the proposed adaptive CCTFD is compared with that of the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs and the adaptive radial Gaussian kernel function-based CCTFD.
In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the $1$-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively. For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively. Additionally, the filtering method utilizing the adaptive radial Gaussian kernel function-based CCTFD is referred to as the adaptive radial Gaussian kernel.
@@
In examples 1–4, additive white Gaussian noise is introduced to the signals, with the SNR ranging from $5\mathrm{dB}$ to $5\mathrm{dB}$ and the observation interval is set to $5\mathrm{s},5\mathrm{s}]$. The sampling frequencies are set as follows: $80\mathrm{Hz}$ for example 1, $100\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for example 3 and $175\mathrm{Hz}$ for example 4.
@@
Figs 1(a) and (b) plot respectively the SNR-MSE (logarithm base $10$) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM signal using seven filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, Wiener filter and the proposed adaptive CCTFD. Figs 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these seven filtering methods. Figs 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these seven filtering methods. Figs 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these seven filtering methods. It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, and Wiener filters under different SNR levels.
@@
Figure 1: Performance evaluation of seven denoising methods for the LFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
@@
Figure 2: Performance evaluation of seven denoising methods for the GELFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
@@
Figure 3: Performance evaluation of seven denoising methods for the QFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
@@
Figure 4: Performance evaluation of seven denoising methods for the SFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
@@
Additionally, we investigate the denoising performance of the seven methods under conditions of different colored noise. The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise. The observation interval is set to $5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4. Table 2 presents the MSE and PSNR values of the output signals through the seven filters under the influence of three types of colored noise. It can be seen that for blue noise, our proposed adaptive CCTFD significantly outperforms the other methods. For pink and red noise, our method is superior to the others for LFM and GELFM signals (except for the PSNR value of red noise in the GELFM signal, which is slightly lower than that of the Wiener filter). For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals. This can be attributed to the spectral characteristics of these signals and the correlation structure of colored noise. QFM and SFM signals have rapidly varying or highly oscillatory instantaneous frequencies, and colored noise introduces frequency-dependent distortions that partially align with these variations. The Wiener filter, derived from second-order statistics under the minimum MSE criterion, can optimally attenuate noise across the entire spectrum by exploiting such correlations. In contrast, the proposed adaptive CCTFD follows the same Wiener filtering principle but performs kernel adaptation in the WVD domain, which may not fully capture the fine-scale structures of rapidly varying signals under colored noise. As a result, although the proposed method excels for signals with smoother frequency variations (e.g., LFM and GELFM) and under white or blue noise, its performance can be slightly inferior to the Wiener filter in these particular scenarios. However, overall, our method consistently outperforms the fixed kernel function filtering methods and the adaptive radial Gaussian kernel function filtering method.
@@
Pink noise Margenau-Hill 0.0016-0.0016 0.1592-0.1592 0.0513-0.0513 0.1319-0.1319
Kirkwood-Rihaczek $0.0676$ 0.0050-0.0050 0.0330-0.0330 0.1200-0.1200
Born-Jordan 0.1582-0.1582 0.3053-0.3053 0.1123-0.1123 0.2134-0.2134
Page $0.0996$ 0.0360-0.0360 0.0201-0.0201 0.0582-0.0582
Adaptive radial Gaussian kernel 0.0451-0.0451 0.4237-0.4237 0.0662-0.0662 0.0510-0.0510
Wiener filter 0.4985-0.4985 0.6911-0.6911 $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ 0.3169-0.3169 0.3369-0.3369
Blue noise Margenau-Hill $0.1437$ 0.4563-0.4563 0.0637-0.0637 0.3730-0.3730
Kirkwood-Rihaczek $0.3580$ 0.3711-0.3711 $0.0063$ 0.2404-0.2404
Born-Jordan 0.1057-0.1057 0.5265-0.5265 0.2016-0.2016 0.9358-0.9358
Page $0.3281$ 0.3355-0.3355 $0.0239$ $0.0682$
Adaptive radial Gaussian kernel 0.0481-0.0481 0.4892-0.4892 0.0671-0.0671 0.0191-0.0191
Wiener filter 0.4676-0.4676 0.6632-0.6632 0.5208-0.5208 1.0831-1.0831
@@
Red noise Margenau-Hill $0.0655$ 0.3004-0.3004 0.0624-0.0624 0.2233-0.2233
Kirkwood-Rihaczek $0.1243$ 0.2567-0.2567 0.0475-0.0475 0.2145-0.2145
Born-Jordan 0.0907-0.0907 0.3810-0.3810 0.1369-0.1369 0.3623-0.3623
Page $0.1720$ 0.2790-0.2790 0.0082-0.0082 0.0851-0.0851
Adaptive radial Gaussian kernel $0.0105$ 0.3455-0.3455 0.0743-0.0743 0.0395-0.0395
Wiener filter 0.8039-0.8039 1.2798-1.2798 $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ 0.4032-0.4032 0.3919-0.3919